### v1

# Zero-Shot Video Editing through Adaptive Sliding Score Distillation

Abstract
The burgeoning field of text-based video generation (T2V) has reignited significant interest in the research of controllable video editing. Although pre-trained T2V-based editing models have achieved efficient editing capabilities, current works are still plagued by two major challenges. Firstly, the inherent limitations of T2V models lead to content inconsistencies and motion discontinuities between frames. Secondly, the notorious issue of over-editing significantly disrupts areas that are intended to remain unaltered. To address these challenges, our work aims to explore a robust video-based editing paradigm based on score distillation. Specifically, we propose an Adaptive Sliding Score Distillation strategy, which not only enhances the stability of T2V supervision but also incorporates both global and local video guidance to mitigate the impact of generation errors. Additionally, we modify the self-attention layers during the editing process to further preserve the key features of the original video. Extensive experiments demonstrate that these strategies enable us to effectively address the aforementioned challenges, achieving superior editing performance compared to existing state-of-the-art methods. For more information, please visit our project homepage: https://nips24videoedit.github.io/zeroshot_videoedit/
11footnotetext: Equal Contribution.22footnotetext: Corresponding Author.
1 Introduction
Video editing, as a task of interactive and controllable generation, has consistently garnered significant attention from the Artificial Intelligence Generated Content (AIGC) community. These techniques enable the modification of existing visual content to meet specific expectations, thereby finding extensive applications in fields such as photography, cinematography, and artistic creation. Despite substantial early explorations, the recent proliferation of Text-to-Image (T2I) and Text-to-Video (T2V) diffusion models has once again spurred a significant amount of related research.
Building on the foundation of these pre-trained diffusion models, existing research can be broadly categorized into Zero-Shot and One-Shot based on their training paradigms. Zero-shot approaches [22, 16, 9] utilize DDIM (Denoising Diffusion Implict Model) Inversion [11] to retain partial information from the original video and incorporate editing guidance as well as temporal priors through attention manipulation and inter-frame attention operations without any training  [5, 14, 31]. In contrast, One-shot approaches [34, 32, 24] attempt to fine-tune pre-trained models to learn subject-independent motion patterns, thereby enabling the editing of video content through prompt replacement. Despite the efficiency and impressive performance of Gaussian noise inversion in diffusion models, this paradigm also introduces two interrelated issues in video editing. Firstly, despite attempts to incorporate DDIM inversion, noise-based inference inevitably disrupts the content and temporal continuity of the source video; Secondly, due to the inherent issues in open-source video generation models, the generated videos frequently exhibit content misalignment and motion flicker, which are even more pronounced to existing video editing works.
A naive solution is to design a more robust update paradigm to replace noise-based ones, which is structure-agnostic and can be seamlessly integrated into most existing works. As in [21] in text-to-3D works, Score Distillation Loss (SDS) is a widely adopted method that aligns with our objective of maintaining consistency. Furthermore, a similar concept Delta Denoising Score (DDS) has also been proven effective in image editing [10], where DDS achieves image editing by directly optimizing the parameters in the source image, ensuring that -edited regions remain largely unaffected. Similarly, this work develops a novel video score distillation paradigm to achieve video-based parameters updating directly. Despite the remarkable success of the score distillation in 3D generation and image editing, experiments have shown that directly applying them [10, 21] with existing open-source text-to-video models [30, 27] yields unsatisfactory results. The edited videos exhibit spotty or blocky bright artifacts, which we refer to as Contaminations. We hypothesize that this issue stems from the weak robustness and unstable generative performance in T2V models, in contrast to T2I models, which adversely affects gradient computation. Undeniably, this issue is exacerbated in video-based update paradigms, as the latent codes of the original video, being at a high resolution and low abstraction level closer to the pixel layer, are particularly sensitive to these errors. Furthermore, image score distillation methods lack video-specific priors and global understanding capabilities, thereby exacerbating the contaminations of the source information.
Figure 1: Our video editing results. More experimental results and complete videos can be found on our project homepage.
This work designs an Adaptive Sliding Score Distillation (ASSD) method, which aims to introduce more robust supervision and video priors to mitigate the aforementioned issues. Specifically, we firstly introduce more stable T2I models to assist in latent code optimization, thereby compensating for the instability of video generation models. To adapt SDS for video data, we propose a local score distillation strategy based on the sliding time-window, which incorporates inter-frame continuity priors during the update process. However, merely relying on local priors is often inefficient, as a significant amount of contaminations in complex video data is difficult to address. Therefore, we design a global gradient-based contaminations prediction strategy to guide the entire distillation process. Additionally, to ensure that the information in the source video is not adversely affected by erroneous generation from the T2V model, specific spatial self-attention layers in the denoising U-Net within the editing branch are modified. Through our proposed weighted attention fusion, we incorporate reference branch information into the gradient computation for optimizing the latent code, correcting errors in the editing branch, and thus enhancing the quality of the update gradients. Extensive experiments on both real and synthetic video editing demonstrate the effectiveness of our proposed method, achieving comparable or superior results to other video editing approaches. In summary, the contributions can be summarized as:
  • We innovatively propose a zero-shot pipeline that addresses video editing by directly optimizing the original video through score distillation.
  • To tackle the contaminations in T2V score distillation, we design ASSD (Adaptive Sliding Score Distillation) to introduce global and local priors for gradient calculation. Additionally, we employ the stable T2I model to provide auxiliary guidance during the editing process.
  • We modify certain spatial self-attention layers in the denoising U-Net, enhancing the robustness of the editing process by preserving information from the original video.
  • Extensive experiments on real-world videos demonstrate that our method achieves comparable or better performance than existing methods in terms of content consistency, motion continuity, adherence to the editing text, and overall video quality.
Figure 2: Overview of Pipeline. Our pipeline includes an upper reference branch and a lower editing branch. We use ASSD for T2V models and DDS for T2I models to jointly derive the gradients for updating the latent code of the video. Additionally, we modify the computation in the spatial self-attention layers of the editing branch to incorporate information from the reference branch.
2 Related Work
Text-to-Video Generation.
Early works on conditional video generation utilized frameworks such as variational autoencoders [15], generative adversarial networks [35, 4, 25, 28], or autoregressive transformers [7, 13, 33, 29], which were trained on small domain-specific datasets. Despite achieving promising results, these approaches struggled to generalize to scenarios beyond the training datasets. Recently, numerous efforts have been made to extend the success of diffusion models in text-to-image generation to the task of text-to-video generation [8, 20]. By training on large-scale image and video datasets, these models have achieved remarkable generation results and controllability over the generated content. To accommodate the input data type of videos, the prevailing methods involve extending the convolutional and attention layers of text-to-image diffusion models along the temporal dimension, such as incorporating temporal convolutions, temporal attention layers, or introducing attention mechanisms between different frames of the video [26, 6, 3, 30, 36]. In this work, our objective is to harness the capabilities of pre-trained text-to-video diffusion models to perform zero-shot editing of original videos, given the corresponding text prompts and edited text prompts. This is achieved solely by utilizing the outputs of the pre-trained model, without subjecting it to any further training.
Text-based Video Editing.
With the success achieved in the field of text-based video generation, many works have attempted to utilize text-based image and video generation models for text-conditioned video editing. Given a reference video with corresponding text prompts and a target text prompt, the video editing task aims to edit the reference video according to the target text prompt.
Tune-A-Video [34] introduces the ability to model temporal information for 2D diffusion models by fine-tuning pre-trained text-to-image diffusion models on reference videos while using DDIM inversion to provide layout and motion guidance from the reference video for editing. Video-P2P [16] adapts the cross-attention control method from the image editing work Prompt-to-Prompt [11] to video editing, employing a dual-branch pipeline that combines the unedited region from the reconstruction branch and the edited region latent features from the editing branch according to the cross-attention map, achieving editing effects in the edited regions while preserving the unedited regions of the reference video. FateZero [22], on the other hand, directly manipulates the intermediate features during the diffusion model generation process under the guidance of the cross-attention map, realizing zero-shot video editing while preserving information from the reference video using DDIM inversion. FLATTEN [5] adopts a similar pipeline to FateZero but introduces guidance from optical flow predicted from the video. TokenFlow [9] and CoDeF [18] embody similar ideas: TokenFlow performs editing on key frames during the generation process and propagates the edits from the key frames to the entire video through feature matching. CoDeF decomposes the video into a canonical content field containing the static content of the entire video and a time-dependent deformation field that includes the transformations needed from the canonical to each frame of the video. By editing the canonical content field and utilizing the deformation field, the editing results are propagated to the entire video.
These works have achieved excellent video editing results, but they typically require executing a complete generation process to obtain the edited video. Although methods such as DDIM inversion have been employed, they still lose some information from the original video, resulting in the problem of maintaining unedited content and motion of the edited subject from the original video. In contrast to their work, our approach directly optimizes the original video by leveraging the guidance provided by pre-trained text-to-video models, better preserving the unedited content from the original video and utilizing the motion information of the edited subject from the original video to achieve stronger inter-frame motion continuity.
3 Method
3.1 Preliminaries
Text-conditioned Latent Video Diffusion Models.
We utilize pre-trained text-conditioned latent video diffusion models [30, 27] to provide guidance for video editing. The text-conditioned latent video diffusion model performs forward noise addition and reverse denoising operations in the latent space of an autoencoder $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$. We denote a video sequence consisting of $N$ frames as $\bm{X}_{0}=\{\bm{x}_{0}^{1},...,\bm{x}_{0}^{N}\}$. During the training phase, the latent input $\bm{Z}_{0}=\{\bm{z}_{0}^{1},...,\bm{z}_{0}^{N}\}=\mathcal{E}(\bm{X}_{0})$ is perturbed to $\bm{Z}_{t}=\{\bm{z}_{t}^{1},...,\bm{z}_{t}^{N}\}$ through the forward diffusion process:
[…]
for $t\in[1,T]$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}1-\beta_{s}$, and $\beta_{s}$ is the variance schedule for timestep $s$. Subsequently, the 3D denoising U-Net is trained to predict the added noise given condition text embedding $y$ using the following objective function:
$\mathcal{L}=\mathbb{E}_{\mathcal{E}(\bm{X}_{0}),y,\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),t}\left[\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,y)\|_{2}^{2}\right],$ (2)
where $\bm{\epsilon}_{\theta}$ represents the 3D denoising U-Net. In text-conditioned generative inference, CFG (Classifier-Free Guidance) [12] is commonly used to strengthen the effect of the condition. The noise prediction in this context is calculated using the following formula:
$\bm{\epsilon}_{\theta}^{w}(\bm{Z}_{t},t,y)=\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,\emptyset)+w(\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,y)-\bm{\epsilon}_{\theta}(\bm{Z}_{t},t,\emptyset)),$ (3)
where $\emptyset$ is null text embedding and $w$ is guidance scale.
Score Distillation for Image Editing.
The SDS loss was first introduced in the text-to-3D work DreamFusion [21], utilizing a pre-trained text-to-image diffusion model to provide update guidance for implicit 3D representations such as NeRF (Neural Radiance Fields) [17]. However, recent work [10] demonstrates that this score distillation loss is equally effective for image editing. Denote the pre-trained T2I model as $\bm{\epsilon}_{\theta}$. Given the parameterized latent code $\bm{z}_{0}(\phi)$ of an image to be edited, the source text embedding $y$, and the target text embedding $y^{*}$, we can use the SDS loss to directly optimize the image to be edited, which can be expressed as:
$\mathcal{L}_{\mathrm{SDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}(\bm{z}_{t}(\phi),t,y^{*})-\bm{\epsilon}\right\|_{2}^{2},\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}),$ (4)
for $t\in[1,T]$. Although the SDS loss provides update guidance for $\bm{z}_{0}(\phi)$ to align with the target text prompt, it often results in over-saturation and blurriness, making the image appear cartoon-like [10]. To address this issue, DDS introduces an additional reference branch aligned with the source text prompt. By calculating the delta score between the model outputs of the two branches, it provides more accurate guidance for updates in the edited parts of $\bm{z}_{0}(\phi)$, while preserving the unedited parts. Let the reference latent code be denoted as $\bm{\hat{z}}_{0}$, which is initialized as the original latent code of the image. The DDS loss can be formulated as:
$\mathcal{L}_{\mathrm{DDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}(\bm{z}_{t}(\phi),t,y^{*})-\bm{\epsilon}_{\theta}^{w}(\bm{\hat{z}}_{t},t,y)\right\|_{2}^{2}.$ (5)
Figure 3: Adaptive Sliding Score Distillation. ASSD adaptively extracts masks for regions with significant gradient changes based on the temporal variation of the update gradients. It then employs a sliding window-based smoothing operation to mitigate errors in the noise predictions from the reference branch.
3.2 Adaptive Sliding Score Distillation
Using a pre-trained text-conditioned latent video diffusion model, we can naturally extend the DDS loss used for image editing to videos. In this work, we use the open-source T2V diffusion models ModelScopeT2V and ZeroScope, denoted as $\bm{\epsilon}_{\theta}$. The video DDS loss can be expressed as:
$\mathcal{L}_{\mathrm{Video-DDS}}(\phi;y^{*})=\left\|\bm{\epsilon}_{\theta}^{w}\left(\bm{Z}_{t}(\phi),t,y^{*}\right)-\bm{\epsilon}_{\theta}^{w}(\bm{\hat{Z}}_{t},t,y)\right\|_{2}^{2},$ (6)
and the gradient used to update $\bm{Z}_{0}(\phi)$ can be calculated as:
$\nabla_{\phi}\mathcal{L}_{\mathrm{Video-DDS}}=\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\frac{\partial\bm{{Z}}}{\partial\phi}.$ (7)
The calculation of this gradient involves a one-step noise prediction from the pre-trained T2V diffusion model. Due to the significantly weaker performance of existing open-source T2V diffusion models in video generation tasks compared to the performance of Stable Diffusion in image generation, this one-step noise prediction contains substantial errors, and becomes particularly pronounced at large timesteps. This leads to unacceptable outcomes when directly transferring the DDS loss to video editing, as evidenced by the results shown in the ablation study in Fig. 5. In the editing branch, the significant variation of the noise predictions at different locations is because the predicted noise also contains information related to the edits. In contrast, the noise predictions in the reference branch should not exhibit such drastic fluctuations. Therefore, we employ a sliding window approach to smooth the noise prediction results of the reference branch along the temporal dimension. The width of the sliding window should not be set too large, as excessive smoothing could lead to blurring of the video. Thus, we selected a small sliding window width of 3. For convenience, we temporarily omit the subscript $\theta$ and superscript $\omega$, let $\bm{\hat{\epsilon}}_{t}^{i}$ denote the noise prediction for the $i$-th frame of the video by the reference branch at timestep $t$. The sliding window-based smoothing operation for this frame is then defined as:
$\operatorname{SS}(\bm{\hat{\epsilon}}_{t}^{i})=\gamma_{i-1}\cdot\bm{\hat{\epsilon}}_{t}^{i-1}+\gamma_{i}\cdot\bm{\hat{\epsilon}}_{t}^{i}+\gamma_{i+1}\cdot\bm{\hat{\epsilon}}_{t}^{i+1},$ (8)
where $\gamma_{i-1},\gamma_{i},\gamma_{i+1}$ are hyper-parameters that control the degree of smoothing, and $\operatorname{SS}(\cdot)$ denotes our sliding window-based smoothing operation. However, applying this sliding window-based smoothing across the entire video can still introduce unnecessary blurring, especially in regions without large prediction errors. To mitigate the impact, guidance is needed to determine where to apply the sliding window-based smoothing. Considering that these prediction errors are typically concentrated around the edited parts, where the update gradients exhibit significant variations along the temporal dimension, we can construct a mask to guide the range of the sliding window-based smoothing operation accordingly. We adaptively obtain the mask through the following process:
$\bm{m}=\left\{\begin{array}[]{ll}1,&\mathrm{abs}\left(\bm{\epsilon}_{t}-\bm{\hat{\epsilon}}_{t}\right)<\mathrm{max}\left(\mathrm{abs}\left(\bm{\epsilon}_{t}-\bm{\hat{\epsilon}}_{t}\right)\right)-\tau\\ 0,&\text{otherwise}\end{array},\right.$ (9)
where $\bm{\hat{\epsilon}}_{t}$ is the noise prediction for all frames output by the reference branch, $\bm{{\epsilon}}_{t}$ is the noise prediction for all frames output by the editing branch, $\operatorname{max}(\cdot)$ denotes the maximization operation along the temporal dimension, and $\tau$ is a hyper-parameter used to control the size of the mask’s effective range. When $\tau$ is smaller, the mask’s effective range is smaller and more stringent, and vice versa. Based on this adaptively extracted mask, the final gradient used to update $\bm{Z}_{0}(\phi)$ is represented as follows:
$\begin{split}\nabla_{\phi}\mathcal{L}_{\mathrm{ASSD}}&=\left(\bm{m}\odot\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\mathrm{SS}\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\right)\right.\\ &\quad+\left.(\bm{1}-\bm{m})\odot\left(\bm{\epsilon}_{\theta}^{\omega}\left(\bm{Z}_{t},y^{*},t\right)-\bm{\epsilon}_{\theta}^{\omega}\left(\bm{\hat{Z}}_{t},y,t\right)\right)\right)\frac{\partial\bm{{Z}}}{\partial\phi}.\end{split}$ (10)
In our proposed adaptive sliding score distillation (ASSD), we mitigate the impact of model prediction errors using a sliding window along the temporal dimension. Additionally, we construct a mask adaptively to minimize the negative effects of this smoothing operation, such as blurring in regions of the video that do not require smoothing. This approach effectively reduces the occurrence of contaminations in the edited video results.
3.3 Weighted Attention Fusion
Although the proposed ASSD effectively reduces errors in the update gradient, local editing errors and changes in unedited parts still occur, which can be observed in Fig. 5. It is desirable to preserve some information in the original video, including the layout of the original video, the details in the unedited parts, and some attributes in the edited parts such as the motion pattern of the edited object. Moreover, some features from the original video can also be leveraged to correct erroneous parts in the edited video to further enhance the editing quality. To better preserve the information and leverage some features in the original video, we introduce a weighted attention fusion.
In the spatial attention layers of ModelScopeT2V or Zeroscope, the self-attention for the latent code $\bm{z}^{i}$ corresponding to the $i$-th frame in the video is calculated as:
$\mathrm{SelfAttn}_{i}=\mathrm{Softmax}(\frac{\bm{W}_{q}(\bm{z}^{i})\bm{W}_{k}(\bm{z}^{i})^{\top}}{\sqrt{d}})\bm{W}_{v}(\bm{z}^{i}),$ (11)
where $\bm{W}_{q}(\cdot)$, $\bm{W}_{k}(\cdot)$, $\bm{W}_{v}(\cdot)$ are linear projections, and $d$ is the scaling factor. We use $\bm{z}^{i}$ to denote the latent code corresponding to the $i$-th frame in the editing branch of ASSD. Correspondingly, we use $\bm{\hat{z}}^{i}$ to denote the latent code corresponding to the $i$-th frame in the reference branch of ASSD. The proposed reference-attention between them is defined as:
$\mathrm{RefAttn}_{i}=\mathrm{Softmax}(\frac{\bm{W}_{q}(\bm{z}^{i})\bm{W}_{k}(\bm{\hat{z}}^{i})^{\top}}{\sqrt{d}})\bm{W}_{v}(\bm{\hat{z}}^{i}).$ (12)
Then, our weighted attention fusion can be implemented as:
$\mathrm{AttnFusion}_{i}=\lambda\cdot\mathrm{SelfAttn}_{i}+(1-\lambda)\cdot\mathrm{RefAttn}_{i},$ (13)
where $\lambda\in[0,1]$. This module allows for the fusion of features from the editing branch with those from the reference branch. We replaced the self-attention calculations in the last few spatial attention layers of the U-Net in the editing branch with our proposed weighted attention fusion, and find this design significantly improves the quality of the edited videos.
3.4 Joint Guidance Incorporating Stable Diffusion
As validated in DDS, the one-step predicted noise provided by Stable Diffusion can offer sufficient quality gradient information for image editing. Therefore, we propose a joint guidance strategy that introduces Stable Diffusion to provide additional gradients for updating $\bm{Z}_{0}(\phi)$. Stable Diffusion uses DDS to separately calculate the gradient for updating each frame in the video. For the $i$-th frame in the video, the gradient can be calculated as follows,
$\nabla_{\phi}\mathcal{L}_{\mathrm{DDS-StableDiffusion}}=\left(\bm{\epsilon}_{\Omega}^{\omega}\left(\bm{z}_{t}^{i},y^{*},t\right)-\bm{\epsilon}_{\Omega}^{\omega}\left(\bm{\hat{z}}_{t}^{i},y,t\right)\right)\frac{\partial{\bm{z}^{i}}}{\partial\phi},$ (14)
where $\bm{\epsilon}_{\Omega}$ denotes a Stable Diffusion V1.5 in our implementation. Finally, the total gradient used to update $\bm{Z}_{0}(\phi)$ can be expressed as:
$\nabla_{\phi}\mathcal{L}_{\mathrm{total}}=\nabla_{\phi}\mathcal{L}_{\mathrm{ASSD}}+\nabla_{\phi}\mathcal{L}_{\mathrm{DDS-StableDiffusion}}.$ (15)
Figure 4: Qualitative Comparison. Comparison of our ASSD with five baseline methods. Our method achieves a good balance between editing and preserving unedited information in the original video. Additionally, it possesses a certain capability for geometric shape editing.
4 Experiments
4.1 Experiments Setting
We collected approximately 30 text-video pairs from public datasets such as DAVIS [19], WebVid [2], and YouTube [1], and extracted 8 or 16 evenly spaced frames from each video for our evaluation data. We used ModelscopeT2V and Zeroscope as our base models. For optimizing the latent code of a video, we employed the SGD optimizer with a learning rate of 0.1, running the optimization for 200 steps, the process takes approximately 3 to 4 minutes on a single V100 GPU.
We evaluated our method alongside five high-performing and representative video editing methods. Tune-A-Video (TAV) [34] fine-tunes the denoising U-Net on a single video to learn the motion patterns of the video to be edited. FateZero [22] achieves zero-shot video editing through DDIM Inversion and attention manipulation techniques. FLATTEN [5] utilizes an external optical flow detection model to enable more accurate attention manipulation. TokenFlow [9] ensures the consistency of the edited video by modifying keyframes and propagating these modifications throughout the entire video. ControlVideo [37] utilizes information such as depth and edges from the original video to control the editing results.
4.2 Qualitative Experiment
Fig. 4 presents the qualitative comparison results between our method and other SOTA baselines. Our method achieves comparable or superior performance in terms of alignment with the target text prompts, preservation of the original video’s layout, motion patterns, and preservation of unedited regions.
Table 1: Quantitative comparisons of our ASSD against SOTA video edit methods.
Methods Auto Metrics User Study
CLIP-Text Align CLIP-Consistency Alignment Consistency Preservation
TAV [34] 0.2988 0.9086 19.06 2.64 4.25
FateZero [22] 0.2738 0.9324 20.63 26.17 30.62
ControlVideo [37] 0.2916 0.9301 8.48 5.30 3.70
FLATTEN [5] 0.2931 0.9249 14.82 17.99 11.11
TokenFlow [9] 0.2910 0.9561 12.67 20.11 24.40
ASSD (Ours) 0.2899 0.9113 24.34 27.79 25.93
4.3 Quantitative Experiment
Auto Metrics.
We utilized the CLIP model [23] for automatic evaluation, employing two metrics: CLIP Text Alignment Score and CLIP Frame Consistency Score. CLIP Text Alignment Score is defined as the mean cosine similarity between the CLIP image embeddings of the frames in the edited video and the CLIP text embedding of the target text prompt. This metric measures how well the edited video aligns with the target text prompt. CLIP Frame Consistency Score is defined as the mean cosine similarity between the CLIP image embeddings of frames within the video. This metric assesses the inter-frame consistency of the edited video. The results of the automatic metric are reported in Tab. 1.
User Study.
Due to the current lack of appropriate evaluation metrics in the field of video generation and editing, automatic assessments based on CLIP often do not align well with the perceptual quality of videos. Therefore, we primarily validate the effectiveness of our method through user studies. We invited 40 volunteers to participate in our user study. The volunteers were asked to evaluate the edited videos based on three criteria: (1) Alignment with the target text prompt. (2) Inter-frame consistency of the edited video. (3) Preservation of the original video’s motion patterns and unedited regions. For each criterion, the volunteers chose the best-performing edited result from those produced by our method and the five baselines. We then calculated the proportion of times each method’s edited results were selected as the best for each criterion. As shown in Tab. 1, our method achieves comparable or superior performance across all three evaluation metrics compared to other SOTA methods.
Figure 5: Abalation of ASSD, Weighted Attention Fusion and Joint Guidance with Stable Diffusion. In most cases, using all the components we propose achieves the best balance between editing and minimizing contaminations.
4.4 Ablation Study
We conduct ablation studies below to validate the effect of each proposed component. As shown in the first row of experimental results in Fig. 5, directly applying DDS for video editing leads to the presence of contaminations near the edited objects in the resulting video. The results in the second row show that by using ASSD instead of DDS, and smoothing the gradient information with adjacent frames, the contaminations in the resulting video are reduced, leading to improved video quality. When we further incorporate the proposed weighted attention fusion mechanism, using features from the reference branch, the contaminations are further reduced. Additionally, we observe that the edited video retains more information from the original video, such as colors and other details. Due to the significantly weaker robustness and generalization to different text conditions of current open-source text-to-video models compared to Stable Diffusion, the latter can sometimes provide higher-quality update gradients for certain text prompts, despite lacking the ability to model temporal information. As shown in the last row of Fig. 5, the updated gradients from Stable Diffusion further reduce some of the contaminations in the video and enhanced the detail quality of the edited subjects.
5 Conclusion and Limitations
We propose a new score distillation-based paradigm for video editing and introduce a novel score distillation method called Adaptive Sliding Score Distillation (ASSD). ASSD mitigates errors in the gradient information provided by text-to-video models through an adaptive sliding window-based smoothing operation and adaptively extracted masks. Our pipeline also incorporates auxiliary guidance from Stable Diffusion to provide more precise gradients for updating the latent code of original video. Additionally, to better capture the information from the original video to correct errors in the edited video and further improve the quality of the update gradients, we employ proposed weighted attention fusion. This modifies certain spatial self-attention layers in pre-trained text-to-video models to support the fusion of information between the editing branch and the reference branch. Experimental results validate the effectiveness and potential of our new paradigm in video editing tasks, as well as the efficacy of our proposed components.
Limitations
It should be noted that the performance of our method heavily relies on the capabilities of the text-to-video models providing the gradient information for editing. The effectiveness of such methods relies on the premise that the pre-trained text-to-video model can provide an update gradient that, despite containing errors, is still effective. Given the current lack of sufficiently powerful open-source text-to-video models, the upper bound of our method’s performance is significantly constrained.